OpenStudy (anonymous):

prove:cos^3 x-sin^3 x = (cosx -sinx)(1+1/2 sin 2x)

OpenStudy (koikkara):

\(cos^3 x-sin^3 x = (cosx -sinx)(1+1/2 sin 2x)\).. do RHS first.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then what

OpenStudy (koikkara):

sorry for delay, i fell into the den of many Q-cats... lol

OpenStudy (anonymous):

its okay @koikkara

zepdrix (zepdrix):

So Koi was suggesting that we apply our `Difference of Cubes` Formula :)\[\large\rm a^3-b^3=(a-b)(a^2+ab+b^2)\]

zepdrix (zepdrix):

Do you see how this matches our formula?\[\large\rm \cos^3 x-\sin^3 x = (\cos x-\sin x)(\cos^2x+\cos x~\sin x+\sin^2x)\]

zepdrix (zepdrix):

Oh he said do right hand side first, I guess he was suggesting something else :)) hehe

zepdrix (zepdrix):

I'm showing the path if we start on left side.

OpenStudy (anonymous):

well i started with the right side because it is more complex

zepdrix (zepdrix):

Hmm I think right hand side is going to be a lot more difficult :d sec thinking.

OpenStudy (anonymous):

okay so what do i do after applying the difference of cubes formula

zepdrix (zepdrix):

Do you notice anything going on right here?\[\large\rm \cos^3 x-\sin^3 x = (\cos x-\sin x)(\color{orangered}{\cos^2x}+\cos x~\sin x\color{orangered}{+\sin^2x})\]It's one of your Pythagorean Identities.

OpenStudy (anonymous):

yes \[\sin^2x \theta+\cos^2 \theta =1\]

zepdrix (zepdrix):

Ok that brings us here,\[\large\rm \cos^3 x-\sin^3 x = (\cos x-\sin x)(\color{orangered}{1}+\cos x~\sin x)\]

zepdrix (zepdrix):

Recall your Double-Angle Formula for Sine? :)

zepdrix (zepdrix):

We'll have to be a little sneaky to make it work for us.

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

\[\sin(2x)_2sinxcosx\]

zepdrix (zepdrix):

Hehe something got a little mixed up there :) I think you meant to write this:\[\large\rm \sin(2x)=2\sin x \cos x\]Let's fiddle around with this identity just a tad bit. Let's divide both sides by 2,\[\large\rm \frac{1}{2}\sin(2x)=\sin x \cos x\]Do you see how that will help us?

OpenStudy (anonymous):

yess

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\frac{1}{2}\sin(2x)=\sin x \cos x}\]And we have this orange thing in our expansion,\[\large\rm \cos^3 x-\sin^3 x = (\cos x-\sin x)(1+\color{orangered}{\cos x\sin x})\]\[\large\rm \cos^3 x-\sin^3 x = (\cos x-\sin x)\left(1+\color{orangered}{\frac{1}{2}\sin(2x)}\right)\]So I guess that's our final step, ya? Yay we did it \c:/

OpenStudy (anonymous):

yay its proved